OpenStudy (anonymous):

2x^2-x + 2=0 use the quad equation

OpenStudy (anonymous):

finals review

OpenStudy (saifoo.khan):

\[\frac 14 (1- i \sqrt {15}),\frac 14 (1+ i \sqrt {15})\]